OpenStudy (anonymous):

Solve by Factoring: 3x2 = -9x What are the solutions for x i got x=0 and x=3 correct or no?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

x=0 and x=-3

OpenStudy (anonymous):

^ that's correct.

OpenStudy (anonymous):

so they dont have to equal zero?

OpenStudy (anonymous):

most or i thought all had to equal zero

OpenStudy (anonymous):

You make it equal zero to find it, yes.

OpenStudy (anonymous):

3x2 =12x 3x2 - 12x = 0 3x(x - 4) = 0 3x = 0 or x - 4 = 0 x = 0 or x = 4 thats an ex. from my book and isnt 3 x -3= 9 im confused

OpenStudy (anonymous):

No, since you don't have a constant in that quadratic. To find the zeroes, just factor out an x. 3x^2 + 9x x(3x+9) x = 0 and 3x + 9 = 0

OpenStudy (anonymous):

Your original question was 3x^2 = -9x For 3x^2 = 12x, it would be: x(3x-12) x = 0 and 3x -12 = 0

OpenStudy (anonymous):

ok in other words i should just plucked them in to see what worked to equal -9 not zero

OpenStudy (anonymous):

You're confusing me, I'm not sure what you mean.